OpenStudy (anonymous):

21. Two sets of values are given below. set 1: 37 , 39 , 35 , 41 , 38 set 2: 41 , 39 , 34 , 39 , 37 Which of the following conclusions can be drawn about the means of these two sets? A. The mean for set 2 is equal to the mean for set 1. B. The mean for set 2 is lower than the mean for set 1. C. The mean for set 2 is higher than the mean for set 1. D. The medians of both sets are equal to the means of both sets.

OpenStudy (anonymous):

@Noemi95

OpenStudy (anonymous):

Well, to find the means you add all the numbers given (each set respectively) and then divide by the number of numbers there is. And The median of a set of numbers is the value for which half the numbers are larger and half are smaller.

OpenStudy (anonymous):

so let's see for the first set, if we add all the numbers we get 190 then we divide by 5 (because there is 5 #'rs in set 1 so we get 190/5=38 and for set 2 we get 190 also, so the mean is 38 also

OpenStudy (anonymous):

so, I would say B is the answer.

OpenStudy (anonymous):

meant A!

OpenStudy (anonymous):

can you also look at these two ?

OpenStudy (anonymous):

yeah, let me take a look

OpenStudy (anonymous):

Kay !

OpenStudy (anonymous):

Let's see, for a 30-60-90 right triangle, the legs are in ratio 1:sqrt3:2, where 2 is the hyp, 1 is short leg, and sqrt3 is long leg. So for 19, hyp=28, then short leg=14 [answer is A]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

for 20 it's D

OpenStudy (anonymous):

want to know why/how?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

to find a shaded section of a circle we use the formula: (degrees of central angle)/360*(pi)*r^2 We know the area of the whole circle which is 169(pi)=530.66, BUT we don't know what the radius is. To find the radius of the circle we use formula: (area/pi) and then find sqrt of it. So, 530.66/3.14=169, and the sqrt of 169=13 Now we have all the info we need to find the area of the section. so it's, 120/360*(pi=3.14)*13^2 solve it and we get 176.89 ( answer D)